# Audio-Enhanced Vision-Language Modeling with Latent Space Broadening for High Quality Data Expansion

### v1

@@ Abstract.
Transformer-based multimodal models are widely used in industrial-scale recommendation, search, and advertising systems for content understanding and relevance ranking. Enhancing labeled training data quality and cross-modal fusion significantly improves model performance, influencing key metrics such as quality view rates and ad revenue. High-quality annotations are crucial for advancing content modeling, yet traditional statistical-based active learning (AL) methods face limitations: they struggle to detect overconfident misclassifications and are less effective in distinguishing semantically similar items in deep neural networks. Additionally, audio information plays an increasing role, especially in short-video platforms like TikTok, yet most pretrained multimodal architectures primarily focus on text and images. While training from scratch across all three modalities is possible, it sacrifices the benefits of leveraging existing pretrained visual-language (VL) and audio models. To address these challenges, we propose kNN-based Latent Space Broadening (LSB) to enhance AL efficiency, achieving an up to 9% recall improvement at 80% precision on TikTok’s proprietary datasets. Additionally, we introduce Vision-Language Modeling with Audio Enhancement (VLMAE), a mid-fusion approach integrating audio into VL models, yielding up to another 9% recall improvement at 80% precision. Our methods are successfully deployed in multiple production systems, leading to significant business gains through online A/B experiments.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
In modern search, recommendation, and advertising systems, multimodal content understanding models play a critical role in evaluating content quality and determining the relevance between search queries and candidate items (Devlin et al., 2019; Hou et al., 2024; Sharma et al., 2024). For example, feed recommendation heavily relies on content understanding models to suppress content that can lead to negative user experiences and promote content aligned with users’ interests. In addition, search advertising systems depend on relevance models to assess the degree of relevance between a given query-ad pair. To enable multimodal information integration across visual, textual, and audio data, multimodal models have become widely adopted, significantly enhancing content understanding capabilities. Improving data quality and modality fusion are two critical directions for further advancing multimodal model performance.
@@
(1) A novel active learning strategy based on Latent Space Broadening (LSB) – By identifying semantically similar hard samples from a set of seed bad cases, we leverage k-Nearest Neighbors (kNN) in the latent space, which is derived from the hidden state embeddings of the multi-modality models. We further introduce a Lookalike Threshold (LT) to quantitatively define embedding similarity, enhancing active learning effectiveness(Chacko et al., 2021).
(2) Vision-Language Modeling with Audio Enhancement (VLMAE) – We compare a genuine approach of audio fusion (Dong et al., 2024) and propose VLMAE, which facilitates the modality crossing between audio and vision-language (VL) information, then further perform supervised fine-tuning incorporating domain-specific knowledge. This approach significantly improves the model’s ability to jointly understand video, text, and audio.
@@ 2. Related Work
Active Learning. The theoretical foundations of active learning algorithms have been extensively studied (Dasgupta, 2005; Balcan et al., 2009; Awasthi et al., 2014; Yan et al., 2017; Beygelzimer et al., 2009). However, statistical algorithms guarantees are often not directly applicable to deep neural networks (DNNs). Due to these limitations, practical active learning applications often rely on heuristic approaches to select examples for labeling. For instance, Tong and Koller (2001) and Sun et al. (2023) propose a margin-based selection strategy, while Settles and Craven (2008) and Shen et al. (2004) suggest combining multiple criteria for natural language processing (NLP) tasks. Culotta and McCallum (2005) explore the use of the least confidence criterion in linear CRF models for sequence prediction tasks. A broader overview of the field can be found in the works of Settles (2010) and Olsson (2009).
@@ 3. Approaches
, where $i=1,2,...,T$, and $H_{f}$ serves as the sequence of visual tokens, facilitating downstream integration with textual tokens.
@@ 3.1. Latent Space Broadening
However, due to the non-linearity nature of $f(h;W)$, it is difficult for the kNN algorithm to explicitly guarantee that $h$ and $h^{\prime}$ are close enough in the latent space. In addition, even though for a certain $h^{\prime}$ we are able to guarantee the $\delta$-vicinity, its actual label $l^{\prime}=\text{idx}(y^{\prime})$ could be different from the label $l=\text{idx}(y)$ of $h$, i.e. under search scenario, two pairs of query-videos could be semantically close, but one is ”$prediction-label\;match$” but the other is ”$prediction-label\;mismatch$”. In this situation, we are not able to expand to as many bad cases as expected.
@@ 3.1.2. Lookalike Threshold (LT)
In this method, taking the latent embedding as input, a binary classification model is trained to distinguish whether there is a mismatch between the prediction of the model $output$ and the human annotation $l$.
Here, we define label $d=1$ as inconsistent between $output$ and $l$ ($\arg\max_{i}\hat{y}\neq\text{idx}(y)$), and $d=0$ as consistent between $output$ and $l$ ($\arg\max_{i}\hat{y}=\text{idx}(y)$); a logistic regression (LR) model is trained with variable $U$ and $b$:
@@ 4.1.1. Evaluation Sets
Search Ads: 12.7k randomly sampled query-ad pairs from the past 30 days TikTok Search Ads impression log. Positive sample rate = $2.6\%$.
Clickbait: 40k positive concentrated sampled videos from the past 60 days TikTok daily published new video log. Positive sample rate = $12\%$.
ANSA: 30k positive concentrated sampled videos from the past 90 days TikTok daily published new video log. Positive sample rate = $26.7\%$.
@@ 4.3.1. Effectiveness of Dataset
We compare the difference between human annotation and model prediction as $multi-class\;loss$ on different datasets listed above.
@@
Table 3. LSB case study for Search Ads: each cell’s first row is seed badcase, and the following three cases are expanded by LSB. It is possible to find semantically similar cases as ”michael scott costume” to ”NICHOLAS CHAVEZ COSTUME”. Similarly, in cases of underestimation, the multiclass probability distribution assigns a probability of 0.8 to a score of 00 (”completely irrelevant”), while manual annotations consistently assign a score of $2$ (”moderately relevant”). This also applies to the underestimation in the second group and the overestimation in the first group.

### v2

@@ Abstract.
Transformer-based multimodal models are widely used in industrial-scale recommendation, search, and advertising systems for content understanding and relevance ranking. Enhancing labeled training data quality and cross-modal fusion significantly improves model performance, influencing key metrics such as quality view rates and ad revenue. High-quality annotations are crucial for advancing content modeling, yet traditional statistical-based active learning (AL) methods face limitations: they struggle to detect overconfident misclassifications and are less effective in distinguishing semantically similar items in deep neural networks. Additionally, audio information plays an increasing role, especially in short-video platforms, yet most pretrained multimodal architectures primarily focus on text and images. While training from scratch across all three modalities is possible, it sacrifices the benefits of leveraging existing pretrained visual-language (VL) and audio models. To address these challenges, we propose kNN-based Latent Space Broadening (LSB) to enhance AL efficiency, achieving an up to 9% recall improvement at 80% precision on proprietary datasets. Additionally, we introduce Vision-Language Modeling with Audio Enhancement (VLMAE), a mid-fusion approach integrating audio into VL models, yielding up to another 9% recall improvement at 80% precision. Our methods are successfully deployed in multiple production systems, leading to significant business gains through online A/B experiments.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3737195ccs: Computing methodologies Artificial intelligenceccs: Information systems Data miningccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
In modern search, recommendation, and advertising systems, multimodal content understanding models play a critical role in evaluating content quality and determining the relevance between search queries and candidate items (Devlin et al., 2019; Hou et al., 2024; Sharma et al., 2024; Wang et al., 2025). For example, feed recommendation heavily relies on content understanding models to suppress content that can lead to negative user experiences and promote content aligned with users’ interests. In addition, search advertising systems depend on relevance models to assess the degree of relevance between a given query-ad pair. To enable multimodal information integration across visual, textual, and audio data, multimodal models have become widely adopted, significantly enhancing content understanding capabilities. Improving data quality and modality fusion are two critical directions for further advancing multimodal model performance.
@@
(1) A novel active learning strategy based on Latent Space Broadening (LSB) – By identifying semantically similar hard samples from a set of seed bad cases, we leverage k-Nearest Neighbors (kNN) in the latent space, which is derived from the hidden state embeddings of the multi-modality models. We further introduce a Lookalike Threshold (LT) to quantitatively define embedding similarity, enhancing active learning effectiveness(Chacko et al., 2021; Zhang and Zhang, 2025).
(2) Vision-Language Modeling with Audio Enhancement (VLMAE) – We compare a genuine approach of audio fusion (Dong et al., 2024; Zhao, 2024) and propose VLMAE, which facilitates the modality crossing between audio and vision-language (VL) information, then further perform supervised fine-tuning incorporating domain-specific knowledge. This approach significantly improves the model’s ability to jointly understand video, text, and audio.
@@ 2. Related Work
Active Learning. The theoretical foundations of active learning algorithms have been extensively studied (Dasgupta, 2005; Balcan et al., 2009; Awasthi et al., 2014; Yan et al., 2017; Beygelzimer et al., 2009; Zhou et al., 2024). However, statistical algorithms guarantees are often not directly applicable to deep neural networks (DNNs). Due to these limitations, practical active learning applications often rely on heuristic approaches to select examples for labeling. For instance, Tong and Koller (2001) and Sun et al. (2023) propose a margin-based selection strategy, while Settles and Craven (2008) and Shen et al. (2004) suggest combining multiple criteria for natural language processing (NLP) tasks. Culotta and McCallum (2005) explore the use of the least confidence criterion in linear CRF models for sequence prediction tasks. A broader overview of the field can be found in the works of Settles (2010) and Olsson (2009).
@@ 3. Approaches
, where i=1,2,,Ti=1,2,...,T, and $H_{f}$ serves as the sequence of visual tokens, facilitating downstream integration with textual tokens.
@@ 3.1. Latent Space Broadening
However, due to the non-linearity nature of $f(h;W)$, it is difficult for the kNN algorithm to explicitly guarantee that $h$ and $h^{\prime}$ are close enough in the latent space. In addition, even though for a certain $h^{\prime}$ we are able to guarantee the $\delta$-vicinity, its actual label $l^{\prime}=\text{idx}(y^{\prime})$ could be different from the label $l=\text{idx}(y)$ of $h$, i.e. under search scenario, two pairs of query-videos could be semantically close, but one is ”predictionlabelmatchprediction-label\;match” but the other is ”predictionlabelmismatchprediction-label\;mismatch”. In this situation, we are not able to expand to as many bad cases as expected.
@@ 3.1.2. Lookalike Threshold (LT)
In this method, taking the latent embedding as input, a binary classification model is trained to distinguish whether there is a mismatch between the prediction of the model outputoutput and the human annotation $l$.
Here, we define label $d=1$ as inconsistent between outputoutput and $l$ ($\arg\max_{i}\hat{y}\neq\text{idx}(y)$), and $d=0$ as consistent between outputoutput and $l$ ($\arg\max_{i}\hat{y}=\text{idx}(y)$); a logistic regression (LR) model is trained with variable $U$ and $b$:
@@ 4.1.1. Evaluation Sets
Search Ads: 12.7k randomly sampled query-ad pairs from the past 30 days Search Ads impression log. Positive sample rate = $2.6\%$.
Clickbait: 40k positive concentrated sampled videos from the past 60 days daily published new video log. Positive sample rate = $12\%$.
ANSA: 30k positive concentrated sampled videos from the past 90 days daily published new video log. Positive sample rate = $26.7\%$.
@@ 4.3.1. Effectiveness of Dataset
We compare the difference between human annotation and model prediction as multiclasslossmulti-class\;loss on different datasets listed above.
@@
Table 3. LSB case study for Search Ads: each cell’s first row is seed badcase, and the following three cases are expanded by LSB. It is possible to find semantically similar cases as ”michael scott costume” to ”NICHOLAS CHAVEZ COSTUME”. Similarly, in cases of underestimation, the multiclass probability distribution assigns a probability of 0.8 to a score of 0 (”completely irrelevant”), while manual annotations consistently assign a score of 22 (”moderately relevant”). This also applies to the underestimation in the second group and the overestimation in the first group.